half a pound of gas undergoes a reversible non flow process at with the pressure remaining constant at 50 psia while the volume increases from 4ft^3 to 8ft^3. find the work done, BTU/lb
1
Expert's answer
2020-03-17T10:02:31-0400

As per the given question,

amount of gas"=0.5 pound=\dfrac{0.5}{2.20}kg"

Pressure=50 psia=344738Pascal

Initial volume ="4ft^3"

Final volume ="8ft^3"

we know that work done "=p_1v_1\ln(\dfrac{V2}{V1})"

"=\dfrac{0.5}{2.20}\times 344738\ln{\dfrac{8}{4}}=0.22\times344738\ln{2}"

"=52569.91J"